OpenStudy (anonymous):

Xavier is riding on a Ferris wheel at the local fair. His height can be modeled by the equation H(t) = 20cosine of the quantity 1 over 15 times t + 30, where H represents the height of the person above the ground in feet at t seconds. Part 1: How far above the ground is Xavier before the ride begins? Part 2: How long does the Ferris wheel take to make one complete revolution? Part 3: Assuming Xavier begins the ride at the top, how far from the ground is the edge of the Ferris wheel, when Xavier's height above the ground reaches a minimum? You must show all work.

OpenStudy (anonymous):

\[h(t)=20\cos \(\frac{ \pi }{ 15 }t)+30\]

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@JFraser

OpenStudy (jfraser):

solve for t=0 for part 1, that's simple enough

OpenStudy (anonymous):

-20+30

OpenStudy (anonymous):

@JFraser

OpenStudy (anonymous):

10

OpenStudy (anonymous):

correct? @JFraser

OpenStudy (jfraser):

why is it -20? doesn't the equation have a positive 20?

OpenStudy (anonymous):

cos=-1?

OpenStudy (jfraser):

if t=0, what's \(\frac{\pi}{15} * 0\)?

OpenStudy (anonymous):

0

OpenStudy (jfraser):

then what's the cosine of zero

OpenStudy (anonymous):

0 haha

OpenStudy (jfraser):

so....

OpenStudy (anonymous):

so 50?

OpenStudy (jfraser):

50

OpenStudy (anonymous):

50 feet before ride begins?

OpenStudy (jfraser):

yes

OpenStudy (anonymous):

ok, move onto part two? :)

OpenStudy (jfraser):

how many degrees make up a full revolution of a circle?

OpenStudy (anonymous):

360

OpenStudy (jfraser):

is there a place in the formula to put an angle?

OpenStudy (anonymous):

i dont think so

OpenStudy (jfraser):

pi is your place for an angle

OpenStudy (anonymous):

insert 360 in for pi?

OpenStudy (jfraser):

yes

OpenStudy (anonymous):

what about t

OpenStudy (anonymous):

20cos(24t)+30

OpenStudy (jfraser):

you're solving for t, because it's asking for time

OpenStudy (anonymous):

can you help me solve?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@phi can you please help me?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@Australopithecus

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@tester97

OpenStudy (anonymous):

@nincompoop

OpenStudy (bradely):

Step by step solution http://www.mathskey.com/question2answer/16927/help-please-thank-you